### v2

# UniCode: Augmenting Evaluation for Code Reasoning

Abstract
Current coding benchmarks often inflate Large Language Model (LLM) capabilities due to static paradigms and data contamination, enabling models to exploit statistical shortcuts rather than genuine reasoning. To address this, we introduce UniCode, a generative evaluation framework that systematically probes LLM limits via: (1) multi-dimensional augmentation transforming seed problems into complex variations to disrupt fixed algorithmic patterns; (2) a highly reliable, automated test generation pipeline for scalable evaluation; and (3) fine-grained metrics for rich error signals. Experiments reveal a 31.2% performance collapse in state-of-the-art models on UniCode, primarily driven by deficiencies in conceptual modeling and scalability reasoning rather than syntactic errors. Furthermore, we uncover a “seed-problem regression” where models revert to memorized seed logic rather than following new specifications, signaling a reliance on shortcuts over reasoning. This work validates UniCode as a robust framework to expose model fragility and foster reasoning-oriented code intelligence.
Machine Learning, ICML
1 Introduction
Developing intelligent systems capable of multi-step reasoning remains a cornerstone of AI research (Wei et al., 2022; Guo et al., 2025; Jaech et al., 2024; Comanici et al., 2025; Li et al., 2022a). Competitive programming has emerged as an ideal testbed for evaluating such capabilities (Li et al., 2022c; El-Kishky et al., 2025), not merely for its rigorous evaluation signals, but because it establishes coding as a general-purpose interface for problem-solving. In this context, writing executable code becomes a universal medium for grounding complex reasoning and formalizing solutions across diverse tasks (Zhu et al., 2025; Quan et al., 2025).
However, current benchmarks suffer from an “evaluation paradox”: while LLMs have nearly achieved saturation on standard coding benchmarks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021), they frequently stumble during real-world interactions (Sergeyuk et al., 2025; Weisz et al., 2025). We attribute this discrepancy to three critical limitations in existing evaluation protocols: 1) data contamination and fixed design patterns, which allow models to exploit statistical shortcuts (Figure 1 (d)); 2) limited scalability due to the high cost of human curation (Jain et al., 2024); and 3) a reliance on static datasets that fail to capture the open-ended nature of real-world software engineering (Fodor, 2025; Stanley et al., 2019).
To address these issues, recent research has explored dataset augmentation via perturbation (Li et al., 2024a; Mirzadeh et al., 2024; Orvalho and Kwiatkowska, 2025). Yet, these approaches predominantly focus on superficial variations, such as variable renaming or surface-level rephrasing, rather than altering the underlying reasoning logic. Consequently, they fail to assess whether a model has truly mastered the algorithmic concepts or is merely recalling specific problem structures. To rigorously evaluate adaptability, there is an urgent need for a systematic augmentation framework that mimics real-world challenges through deep, structural transformations rather than shallow manipulations.
In this work, we introduce UniCode, a framework for the Augmented Evaluation of code reasoning (Figure 1 (a)), which employs a generative approach to systematically stress-test models under meaningful structural, compositional, and conceptual shifts. Our framework offers the following three key contributions:
Systematic Task Augmentation. We propose a rigorous methodology that transforms high-quality “seed problems” into a diverse array of challenging tasks in Section 2. Specifically, Atomic variations modify targeted task facets (e.g., narrative, rules, or input scale) to determine if models can adapt to structural changes rather than relying on rote retrieval; Compositional variations integrate multiple knowledge points to generate novel combinatorial patterns, forcing models to engage in genuine multi-step reasoning (see  Figure 1 (b)). By applying broadly meaningful transformations ranging from narrative rewriting to logical reconfiguration, we simulate the complexity of real-world requirements.
Scalable and Robust Evaluation. To address the scalability bottleneck of human curation, we develop a stress-driven synthesis pipeline in Section 3. By integrating stress-testing with multi-model consensus validation, our approach achieves 94.5% reliability in test case generation without requiring ground-truth solutions. Test cases from three distinct sources: random sampling, adversarial boundary cases, and LLM-synthesized hard examples, guarantee a robust assessment of functional correctness and corner-case stability. This allows for the dynamic expansion of an effectively infinite, contamination-resistant problem space, ensuring the benchmark remains challenging as models evolve.
Fine-grained Diagnostic Metrics. Current evaluations often rely on a single 0-1 pass rate, which obscures the specific deficiencies of the model. UniCode transcends this limitation by providing a comprehensive diagnostic toolkit. We categorize errors into modeling flaws, logic gaps, algorithmic complexity misjudgments, and implementation bugs in Section 5.2. This granular analysis moves beyond simple accuracy to reveal why a model fails, offering practical insights for improving code reasoning capabilities.
We benchmark 19 LLMs on UniCode, finding it highly challenging and discriminative. Even the top-performing model, o4-mini, achieves a pass rate of only 70.3%. Our in-depth analysis reveals a critical vulnerability: while LLMs remain robust against superficial perturbations, their performance collapses ($30\%$) under our structural augmentations. Case studies reveal that models often remain “trapped” in the reasoning paradigms of the seed problems, failing to adapt to the augmented logic (Figure 5). By providing a fine-grained view of these failures, UniCode paves the way for the development of more robust, generalized multi-step code agents.
Figure 1: An Overview of UniCode. (a) UniCode Framework. We propose a generative pipeline to create a vast evaluation space via multi-dimensional augmentations to probe reasoning axes, a test generation module for scalability, and fine-grained diagnostics that expose specific failure modes for a transparent assessment of code reasoning. (b) An Example: Probing Code Reasoning Ability via Five Augmentation Axes. We design Atomic Augmentations (narrative, rule, and efficiency) to alter task facets, ranging from surface-level shifts to reconfigurations of the reasoning graph topology; and Compositional Augmentations (sequential composition and concept fusion) to disrupt fixed algorithmic patterns and test combinatorial generalization. (c) Model Fragility: Vulnerability to Logic Alteration. While LLMs are robust under surface-level narrative shifts, performance drops significantly when core logic is modified. The most severe failures in sequential and fusion tasks indicate struggles with long causal chains and algorithmic integration. (d) Beyond Human Curation: Identifying Reasoning Blind Spots. Unlike static benchmarks (e.g., TACO (Li et al., )) with predictable patterns, UniCode explores a broader spectrum of algorithmic combinations to expose “reasoning blind spots” that human designers typically fail to cover. (e) Performance Collapse: Memorization vs. Adaptation. Large-scale testing reveals an average drop of $30\%$ when moving from seed problems to their augmented variations across LLMs. We observe a “seed-problem regression” where models default to the memorized logic of seed problems rather than adapting to new, augmented constraints. Best viewed when zoom in.
Figure 2: Stress-driven pipeline for ground-truth generation. (1) Stress Testing: we use a brute-force solver to produce trusted outputs on small inputs, which serve as stress tests to filter LLM-generated optimized solvers. (2) Consensus Validation: The remaining solvers are executed on large inputs, and the output is selected by strict majority vote. (3) LLM Adjudication: A powerful LLM adjudicates between conflicting outputs for inputs where no majority is reached. The ablation study in App. B.1 verifies the effectiveness of each stage.
2 Augmentation Axes for Code Reasoning
Data-driven LLMs often rely on statistical correlations over logical reasoning, faltering in complex, open-world scenarios (McCoy et al., 2019). While perturbations are standard for testing, existing benchmarks struggle to challenge increasingly robust models.To address this, we design Atomic Augmentations (narrative, rule, and efficiency) ranging from surface shifts to structural reasoning graph reconfigurations; and Compositional Augmentations (sequential composition and concept fusion) to disrupt fixed algorithmic patterns and evaluate combinatorial generalization. These methods are illustrated in Figure 1 (b) and detailed below:
Narrative Perturbation
This axis modifies variable names, thematic backgrounds, or injects irrelevant contextual noise without altering the underlying logic. It specifically probes whether a model suffers from “token bias” (Jiang et al., 2024) or exhibits content-agnostic reasoning. For example, we reframe the abstract Longest Increasing Subsequence (LIS) problem into a real-world scenario like Identifying the Longest Growth Period in Stock Trends (see Figure 1 (b)). If a model fails due to this narrative shift, it indicates a reliance on near-neighbor matching (Li et al., 2024b) of familiar problem descriptions rather than a robust understanding of the logical core.
Rule Modification
Standard programming problems often have “canonical” solutions that LLMs easily memorize. By subtly altering operational rules or boundary conditions, we invalidate these memorized paths (Dziri et al., 2023). A representative transformation is shifting LIS to the Longest Non-Decreasing Subsequence. While seemingly minor, this change shifts the comparison operator (from >> to $\geq$), requiring the model to re-calibrate its logical flow according to novel instructions rather than retrieving pre-trained code snippets, effectively distinguishing retrieval from reasoning.
Efficiency Scaling
Genuine reasoning entails an awareness of the computational budget and the ability to adapt as data scale increases (Zubić et al., 2025). This axis tests if a model can transition from a naive approach to a more optimized algorithm when complexity demands it. For instance, when the input size $n$ for LIS scales to $10^{5}$, the standard $O(n^{2})$ approach becomes computationally prohibitive. The model must recognize this bottleneck and pivot to a greedy strategy ($O(n\log n)$). This transition probes the model’s capacity for high-level strategy selection and its understanding of algorithmic efficiency beyond simple template filling.
Sequential Composition
This dimension involves chaining multiple distinct algorithmic steps to examine the stability of the reasoning chain. As the sequence of operations grows, the probability of failure increases—a phenomenon known as error propagation (Schaeffer et al., 2023). In a composite variant like the Longest Bitonic Subsequence, the model must compute the LIS from both the prefix and suffix and then integrate the results. Such tasks reveal the fragility of the reasoning process, as minor logical flaws that might be hidden in single-step tasks are amplified during intermediate state transfers.
Concept Fusion
Real-world challenges often lie at the intersection of disparate domains. The fusion variant merges distinct algorithmic concepts into a single problem, creating novel combination patterns that human designers often overlook. For instance, while dynamic programming is frequently paired with string manipulation in standard datasets, it is rarely combined with game theory or greedy algorithms, making such intersections particularly difficult to navigate. By creating these original pairings, we probe whether the model can genuinely integrate separate concepts to achieve combinatorial generalization, which is recognized as a cornerstone of human-like reasoning (Battaglia et al., 2018).
By evaluating performance across these strategic axes that reconfigure the reasoning graph topology, we systematically probe the model’s structural generalization and its adaptability to novel coding challenges.
3 Scalable and Rigorous Test Generation
To scale evaluation effectively, an automated test generation pipeline is essential. However, the core challenge is ensuring test quality for novel problems lacking human-authored solutions. We address this by using three distinct input types to cover boundary conditions and attacks, then establishing trusted outputs through a stress-driven pipeline.
3.1 Input Generation
Low-quality test cases often lead to model mis-ranking (Jain et al., 2024; Liu et al., 2025; Wang et al., 2025). To ensure robustness, we construct test cases by prompting LLMs to generate inputs from three complementary sources:
  • Random Generation ($G_{\mathrm{rand}}$): targets general correctness that samples broadly from the valid input space.
  • Adversarial Generation ($G_{\mathrm{adv}}$): targets algorithmic inefficiencies or edge-case failures, such as boundary-value, extreme sequence lengths, and oscillating patterns.
  • Corner Generation ($G_{\mathrm{corn}}$): targets subtle failure modes using challenging and small-scale inputs.
All candidate inputs from each source $\tau\in\{\text{rand, adv, corn}\}$ (input prompts detailed in App. Section B.7) are filtered by a verifier $V$. Let $G_{\tau}$ represent the initial candidate pool; the final set of verified inputs $I_{\tau}$ is defined as:
$I_{\tau}=\{\,x\in G_{\tau}\mid V(x)=\mathrm{true}\,\}.$
To balance coverage and efficiency (Liu et al., 2023a), we specifically assemble a test suite $S$ of diverse 50 cases: 20 random, 20 adversarial, and 10 corner cases (empirically tuned; see App. Section B.2). Table 3 demonstrates that this configuration ensures both correctness and coverage.
Model Difficulty (Pass@1) Test Impact ($\Delta\downarrow$) Avg. Pass@1 Cost / Prob. (\$)
Easy Medium Hard $\Delta_{Rand}$ $\Delta_{Adv}$ $\Delta_{Corn}$
Reasoning Models
o4-mini (high)* 94.9% 78.2% 21.6% $\mathit{-2.5\%}$ $\mathit{-6.6\%}$ $\mathit{-10.2\%}$ 70.3% 0.0269
gpt-5 (medium)* 89.5% 77.6% 18.8% $\mathit{-2.3\%}$ $\mathit{-5.6\%}$ $\mathit{-8.9\%}$ 67.7% 0.0390
o4-mini (medium)* 89.2% 73.6% 20.3% $\mathit{-2.7\%}$ $\mathit{-8.8\%}$ $\mathit{-11.8\%}$ 66.1% 0.0205
google/gemini-2.5-pro* 94.0% 53.1% 8.5% $\mathit{-2.8\%}$ $\mathit{-7.2\%}$ $\mathit{-9.3\%}$ 61.6% 0.2015
deepseek-v3.1 (thinking) 89.2% 59.8% 11.5% $\mathit{-2.0\%}$ $\mathit{-4.1\%}$ $\mathit{-8.1\%}$ 60.5% 0.0276
deepseek-r1 80.3% 36.4% 5.1% $\mathit{-3.2\%}$ $\mathit{-6.8\%}$ $\mathit{-7.3\%}$ 55.6% 0.0250
o3-mini (medium)* 86.2% 50.0% 6.0% $\mathit{-2.8\%}$ $\mathit{-7.8\%}$ $\mathit{-7.0\%}$ 55.1% 0.0230
qwen3-235b-a22b 80.2% 39.7% 5.1% $\mathit{-2.3\%}$ $\mathit{-9.9\%}$ $\mathit{-13.3\%}$ 53.5% 0.0343
gemini-2.5-flash* 81.4% 22.6% 4.8% $\mathit{-2.6\%}$ $\mathit{-7.7\%}$ $\mathit{-6.3\%}$ 47.7% 0.0090
grok-3-mini* 77.8% 21.7% 3.3% $\mathit{-2.1\%}$ $\mathit{-10.0\%}$ $\mathit{-6.4\%}$ 46.4% 0.0035
claude-3.7-sonnet* 76.2% 24.1% 2.4% $\mathit{-1.2\%}$ $\mathit{-11.5\%}$ $\mathit{-5.2\%}$ 45.5% 0.1282
Non-Reasoning Models
deepseek-chat-v3.1 82.7% 29.3% 3.9% $\mathit{-2.0\%}$ $\mathit{-5.1\%}$ $\mathit{-7.3\%}$ 49.8% 0.0068
gpt-4.1-mini* 73.7% 20.9% 3.8% $\mathit{-2.8\%}$ $\mathit{-8.8\%}$ $\mathit{-7.4\%}$ 42.4% 0.0070
gpt-4.1* 62.1% 21.8% 1.4% $\mathit{-3.0\%}$ $\mathit{-10.2\%}$ $\mathit{-10.4\%}$ 36.5% 0.0071
qwen3-coder 66.5% 9.3% 0.0% $\mathit{-1.3\%}$ $\mathit{-10.5\%}$ $\mathit{-9.9\%}$ 35.4% 0.0145
claude-sonnet-4* 60.7% 14.0% 2.0% $\mathit{-1.5\%}$ $\mathit{-13.1\%}$ $\mathit{-5.1\%}$ 32.4% 0.0211
llama-4-maverick 51.3% 8.6% 0.0% $\mathit{-1.3\%}$ $\mathit{-9.5\%}$ $\mathit{-8.8\%}$ 26.2% 0.0006
gpt-4o* 31.3% 2.2% 0.0% $\mathit{-0.4\%}$ $\mathit{-2.2\%}$ $\mathit{-6.4\%}$ 15.4% 0.0139
qwen-2.5-32b-coder 27.2% 2.2% 0.0% $\mathit{-1.8\%}$ $\mathit{-0.3\%}$ $\mathit{-5.5\%}$ 13.4% 0.0038
gemma-3-27b-it 26.1% 2.2% 0.0% $\mathit{-0.2\%}$ $\mathit{-4.3\%}$ $\mathit{-3.5\%}$ 13.1% -
llama-3.3-8b-instruct 11.2% 1.1% 0.0% $\mathit{-0.4\%}$ $\mathit{-0.2\%}$ $\mathit{-0.3\%}$ 5.5% 0.0002
Table 1: UniCode Leaderboard. We report Pass@1 across three difficulty levels and evaluate model robustness via $\Delta$ ($\Delta=\text{Overall}-\text{w/o Test}$), where larger drops indicate greater vulnerability. The significant drops ($\Delta\downarrow$) demonstrate the UniCode’s capability to expose reasoning flaws. The cost per problem is included to help identify cost-effective models for future research. (*: closed models).
3.2 Ground-Truth Construction
Establishing ground-truth outputs for novel problems is challenging. We devise a multi-stage pipeline (Figure 2) that mirrors a rigorous human validation process.
Stage 1: Brute-Force & Solver Filtration We generate a brute-force solver $B$ via LLM to create ground-truth pairs for small-scale inputs ($I_{s}$). Multiple candidates and consensus ensure $B$’s reliability. We then prompt LLMs for $M$ optimized candidate solutions $\{C_{1},\dots,C_{M}\}$. A candidate enters the trusted pool $\mathcal{P}$ only if it matches $B$ on all $I_{s}$:
$\mathcal{P}=\{C_{j}\mid C_{j}(i)=B(i)\;\forall i\in I_{s}\}.$
Stage 2: Consensus on Large-Scale Inputs For large-scale inputs ($I_{\ell}$) where brute-force is infeasible, we use the pool $\mathcal{P}$. The ground truth for $i\in I_{\ell}$ is determined by a strict majority vote ($\lfloor N/2\rfloor$) among the $N$ optimized solvers.
Stage 3: LLM Adjudication If no majority exists, the top two outputs ($o_{1},o_{2}$) are sent to a high-reasoning LLM (e.g., o4-mini) for analysis. If the LLM yields a decisive judgment, that output is accepted; otherwise, the input is discarded to ensure data integrity. We validate that each component improves test case accuracy in Table 3.
Figure 3: Model performance across reasoning variants. Red dashed lines denote seed problem performance. Key observations: (1) All models exhibit performance variance (e.g., claude-sonnet-4 shows a gap of $60\%$), suggesting substantial disparities in reasoning capabilities across different dimensions. (2) Even reasoning-optimized models (o4-mini, deepseek-r1) are vulnerable to sequential integration, concept fusion and efficiency scaling, which exposes a critical fragility of multi-step reasoning under novel constraints.
4 Benchmark Curation and Leaderboard
This section describes the construction and validation of the UniCode benchmark. We first present our data curation pipeline and a human study evaluating problem quality, followed by a comprehensive leaderboard overview.
4.1 Data Pipeline and Quality
Problem Generation
We curated a dataset of $25,000$ seed problems aggregated from various platforms (e.g., LeetCode, CodeForces). To ensure competitive quality, we removed non-competitive tasks, duplicates, and unclear specifications. Tags were assigned hierarchically (e.g., graph $\rightarrow$ shortest-paths, flow) by an LLM identifying 1–3 core skills per problem (see App. A.6 and B.8). To minimize potential generator bias, as illustrated in App. B.3, we used o4-mini and deepseek-r1 to generate 492 candidate problems following the methods in §2. We excluded low-difficulty problems that were solved perfectly by all baseline models.
Test Suites and Constraints
Each problem includes five components: description $D$, tag set $\mathcal{T}$, time limit $TL$, memory limit $ML$, and test cases $\mathcal{C}$. The time limit $TL$ and memory limit $ML$ are determined by running validated, optimized solutions $\mathcal{O}_{\text{valid}}$:
$\mathrm{TL}=\Big\lceil k\cdot\min_{o\in\mathcal{O}_{\text{valid}}}T(o)\Big\rceil,\qquad\mathrm{ML}=\Big\lceil k\cdot\mathrm{Mem}\big(o^{\star}\big)\Big\rceil,$
where $o^{\star}\;=\;\arg\min_{o\in\mathcal{O}_{\text{valid}}}T(o).$ We select the minimum runtime across validated reference solutions to avoid loose time limits and multiply it by a conservative safety factor $k=3$ to accommodate variations in alternative correct implementations. We then execute $T(\cdot)$ and $\mathrm{Mem}(\cdot)$ within a secure sandbox environment (Bytedance-seed et al., 2025).
Test Suite Quality
To evaluate the quality of the generated problem variants, we assessed 113 sampled problems via a blinded study with five expert annotators. The benchmark achieves a 98.2% validity rate and 92.3% inter-annotator agreement, confirming technical clarity (See App. B.4 for detail). To evaluate the test cases generated by the stress-driven pipeline, we assess correctness (valid solutions accepted) and coverage (invalid solutions rejected). Our ablation study in App. B.1, using Test-Eval (Yang et al., 2025b), demonstrates the effectiveness of our pipeline: reaching 94.5% correctness (vs. 86.9% baseline) and 86.0% coverage (vs. 80.2% baseline), as shown in Table 3. While automatic generation is not error-free, our mathematical proof in App. B.5 confirms that the generated benchmark remains statistically reliable for evaluation.
Release Artifacts.
We will release problem statements, test suites, and metadata (tags, generators, and prompts) to support reproducibility and downstream analysis.
4.2 UniCode Leaderboard
To offer a macroscopic perspective on the UniCode landscape, we evaluate 19 state-of-the-art LLMs across various architectures, parameter scales, and reasoning capabilities.
As shown in Table 1, UniCode is both highly challenging and discriminative, with overall pass@1 scores ranging from 70.3% (o4-mini-high) to 5.5% (llama-3.3-8b-instruct). Performance collapses on the hard split, where several models record 0.0%, underscoring the benchmark’s difficulty. Reasoning-oriented models lead the rankings, validating the effectiveness of test-time compute scaling for complex logical inference. Our results show over 90% alignment with uncontaminated benchmarks (App. A.4), confirming that UniCode is robust against data contamination and provides an unbiased assessment of model performance.
Models generally struggle more with adversarial and corner cases than random generation. For instance, qwen3-235b-a22b show pronounced sensitivity, with $\Delta_{\mathrm{Corn}}$ reaching $13.3\%$. Conversely, deepseek-v3.1 (thinking) exhibits superior stability with minimal performance drops. Furthermore, cost–performance analysis identifies o4-mini (high) as highly efficient, achieving top-tier pass@1 at \$$0.0269$ per problem - nearly $7.5\times$ more cost-effective than gemini-2.5-pro for similar accuracy. While these results reveal UniCode’s difficulty, they do not pinpoint specific failure modes. The following section decomposes performance across five reasoning axes to uncover where and why models fail.
5 In-Depth Code Reasoning Analysis
In this section, we investigate the fragility of LLM code reasoning across diverse variants and introduce a fine-grained taxonomy to categorize the origins of these failures.
5.1 How Fragile is Code Reasoning in LLMs?
To better understand the reasoning ability of LLMs, we curated 132 new problems using livecodebench v1  [^1] Initial Release: May 2023 – March 2024. seed tasks, where models previously excelled. We evaluated a selection of reasoning-specialized, general-purpose, and open-source models across various performance tiers.
General Performance Drop and High Variance
As shown in Figure 3, all models suffered significant performance declines, revealing their fragility when handling structural and conceptual shifts in reasoning tasks. We observe that even state-of-the-art LLMs fail to achieve comprehensive mastery across diverse reasoning paradigms. Furthermore, models exhibit non-negligible variance across test sets; for instance, claude-3.5-sonnet shows a performance gap exceeding 60% between scenarios. This imbalance suggests that aggregate scores often mask significant deficiencies in generalization.
Failure Under Structural Reasoning Alterations
While some LLMs remain robust against narrative perturbations, all models suffer sharp performance declines when the underlying reasoning graph topology is altered. Constraint modifications consistently degrade performance, highlighting the difficulty of preserving logical coherence under novel requirements. Efficiency scaling is particularly challenging: even the specialized o4-mini experiences a 37% drop, marking its weakest dimension. The most significant failures occur in sequential reasoning and concept fusion, indicating persistent difficulty in composing multiple logical components and maintaining long causal chains. Notably, gemma-3-27b-it scored a negligible 1.3% in sequential tasks, representing a near-total loss of functional capacity. Model performance exhibits a continuous declining trend as the augmentation depth increases (see App. A.5).
Seed-problem Regression Phenomenon.
We observe that models often default to original seed-problem logic rather than reasoning from updated task specifications (see Figure 5). This behavior suggests a reliance on heuristic shortcuts rather than rigorous logical deduction. For example, a model may employ a simple parity count because it recognizes a palindrome sub-task, yet fail to integrate new constraints. Alternatively, models rely on “low-efficiency templates” suitable for base problems without accounting for increased input scales. Future research should focus on enhancing models’ zero-shot adaptation to novel constraints.
Figure 4: Error distributions across LLMs. Modeling and complexity errors are the primary failure modes, revealing inherent algorithmic weaknesses. (See Figure 5 for the detailed case study.)
Figure 5: Case study of error patterns in code reasoning. Model failures are not stochastic but exhibit a seed-problem regression. Even when new problem invalidate original P1/P2 logic, LLMs frequently revert to memorized modeling paradigms or suboptimal complexities that were sufficient for the base problems. This suggests that LLMs often rely on heuristic shortcuts rather than true logical synthesis.
5.2 Where Do LLMs Fail in Code Reasoning?
To provide a more comprehensive overview of why LLMs fail, we introduce a systematic taxonomy to analyze their root causes. Modeling error: selecting an incorrect algorithmic paradigm for the task (e.g., opting for dynamic programming when a greedy approach suffices). Logic bug: implementation structural flaws (e.g., incorrect conditional or improper variable handling). Indexing/caching bug: incorrect cache sizing, array bounds violations, or failed boundary checks. Complexity error: utilizing suboptimal algorithms when constraints demand more efficient solutions. Others: minor implementation oversights, including output formatting errors or library usage. We prompt LLMs to categorize errors by providing failed code, passed solutions and failed test cases (see App. Section B.9).
Through a fine-grained analysis of failed cases, we observe an “overhead” phenomenon: models like gemma-3-27b-it and gpt-4o-mini rarely fail due to isolated mistakes. Instead, they exhibit cascading failures where multiple errors occur simultaneously. This indicates that when a task exceeds a certain complexity threshold, the model’s logical coherence may break down, resulting in repetitive, nonsensical code snippets or hallucinatory logic.
As shown in (Figure 4), indexing and others errors account for the smallest fraction of failures. Although dominant error types vary, modeling errors remain a primary challenge, indicating persistent difficulty in problem conceptualization and algorithm selection. Another major contributor is complexity error: models such as gemini-2.5-flash and gemma-3-27b-it struggle with time complexity analysis and appear insensitive to resource constraints. These results suggest that while LLMs are proficient at syntax, they lack a robust grasp of algorithmic efficiency and conceptual modeling.
6 Related Work
Competitive Coding
LLM code generation evaluation is a rapidly evolving field (Jaech et al., 2024; Li et al., 2023; Guo et al., 2024; Hui et al., 2024; Zhang et al., 2023; Guo et al., 2025; Li et al., 2022b; Shao et al., 2024; Allal et al., 2023; Zhao et al., 2025), as code is increasingly viewed as a potential source of reasoning ability (Fu et al., 2022; Li et al., 2022a). Foundational benchmarks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021) use standalone patterns, but their static nature risks data contamination and statistical shortcuts (Oren et al., 2023; Golchin and Surdeanu, 2023; Riddell et al., 2024; Roberts et al., 2023; Tang et al., 2024). Recent benchmarks (Li et al., ; Gu et al., ; Zhu et al., 2025; Chambon et al., ) introduce complex competitive problems to test algorithmic reasoning, yet updates remain slow and patterns fixed (Zheng et al., 2025b; Jain et al., 2024). We study open-ended tasks (Stanley and Lehman, 2015; Clune, 2019; Lehman and Stanley, 2011) explored in recent benchmarks (Xie et al., 2024; Heesch et al., 2025; Chen et al., 2025), which have limited difficulty and reasoning. UniCode generates infinite competitive programming problems via deep structural shifts to probe true code reasoning.
Automated Test Generation
Accurate algorithmic assessment requires scalable test generation, which often relies on manual curation (Chen et al., 2021; Hendrycks et al., 2021; Austin et al., 2021; Li et al., ; Quan et al., 2025). While some use LLMs for test synthesis (Chen et al., 2022; Schäfer et al., 2023; Liu et al., 2023a; Wang et al., 2025; Jain et al., 2024), they frequently depend on existing reference solutions (Schäfer et al., 2023; Tufano et al., 2022; Chen et al., 2022; Liu et al., 2023b). Other works (Chou et al., 2025; Lops et al., 2025; Anand et al., 2013; Sofokleous and Andreou, 2008; Tufano et al., 2020; Swain et al., 2012) generates problems from code but lacks algorithmic depth. UniCode introduces a reliable, stress-driven pipeline enabling robust evaluation for novel problems without ground-truth code.
Generative and Augmented Evaluation
To combat benchmark stagnation (Oren et al., 2023; Roberts et al., 2023), generative evaluation creates dynamic tasks, as seen in game environments (Zheng et al., 2025a), medical diagnosis (Lin et al., 2025), logical reasoning (Parmar et al., 2024), synthetic data structures (Zhu et al., 2023), and agentic tasks (Shi et al., 2025). We adapt this to algorithmic reasoning, using various perturbations and structural variations that may change the underlying reasoning graph (Pei et al., 2025; Huang et al., 2025; Wu et al., 2021) to expose reasoning fragility. Unlike traditional instruction-tuning or Self-Instruct (e.g.,  (Xu et al., 2024; Luo et al., 2023)), or recent robustness benchmarks like (Stolfo et al., 2023; Mirzadeh et al., 2024; Li et al., 2024a; Wang and Zhao, 2024; Yang et al., 2025a; Ramezanali et al., 2025; Orvalho and Kwiatkowska, 2025; Patel et al., 2024) which focus on surface-level perturbations or the linear extension of reasoning steps, UniCode targets underlying algorithmic transformations to disrupt memorized logic and evaluate true generalization.
7 Conclusion
In this paper, we introduce UniCode, a comprehensive and novel generative framework designed to address the “evaluation paradox” in current Large Language Model (LLM) coding benchmarks. To disrupt models’ reliance on statistical shortcuts, we implement multi-dimensional augmentations targeting structural, compositional, and conceptual shifts, supported by a stress-driven synthesis pipeline that ensures a scalable and contamination-resistant evaluation. Our findings reveal a significant 31.2% performance collapse across state-of-the-art LLMs, exposing a critical gap where high scores on static benchmarks fail to translate into robust logical adaptability. Ultimately, this research highlights the fragility of current models under novel logic patterns and underscores the urgent need for reasoning-oriented development in AI coding agents to bridge the gap between benchmark performance and real-world applicability.
Ethics Statement
This paper introduces UniCode, a framework designed to advance the field of machine learning by fostering genuine reasoning-oriented code intelligence. By exposing the “evaluation paradox” where models rely on statistical shortcuts, our work provides a critical foundation for developing more robust, generalizable AI agents. Beyond technical evaluation, UniCode contributes to the broader goal of building reliable and fair AI systems by systematically identifying logic fragilities. Ultimately, this research redefines benchmarking standards and facilitates the creation of safer AI technologies capable of handling complex, real-world reasoning tasks with higher fidelity and transparency.
Appendix A Extended Experimental Results and Diagnostics
A.1 UniCode Efficiency Leaderboard
To illustrate the performance-cost efficiency frontier and assist users in selecting models that balance budget with performance, Figure 6 presents the UniCode leaderboard. This plot maps the Pass@1 score against the average cost per problem for various models, including: gpt-5-2025-08-07, o4-mini-2025-04-16 high, o4-mini-2025-04-16 medium, gemini-2.5-pro, deepseek-v3.1-thinking, deepseek-r1-0528, o3-mini-2025-01-31, qwen3-235b-a22b, gemini-2.5-flash, grok-3-mini, claude-3.7-sonnet:thinking, deepseek-chat-v3.1, gpt-4.1-mini-2025-04-14, gpt-4.1-2025-04-14, qwen3-coder, claude-sonnet-4-20250514, llama-4-maverick:free, gpt-4o-2024-11-20, qwen-2.5-32b-coder, and llama-3.3-8b-instruct.
Figure 6: UniCode Leaderboard: Performance vs. cost efficiency across various models.
A.2 Performance Across Algorithmic Paradigms
To better understand model capabilities, we sample representative models and categorize the problems by their primary algorithmic paradigm. Performance is evaluated using the pass@1 metric. As shown in Figure 7, the results reveal distinct strengths and weaknesses across various problem types.
Figure 7: Comparative performance across problem types for closed-source (left) and open-source (right) models. The x-axis represents algorithmic tags. The y-axis refers to the pass@1 rate. Best viewed zoomed in.
The models demonstrate high proficiency in deterministic, template-driven tasks such as fundamentals and data structures. These problems, which often involve standard data structure manipulations, are likely well-represented in training corpora from textbooks and online repositories. Their solutions typically follow predictable patterns that models can easily recognize and reproduce. In contrast, performance drops on problems requiring novel reasoning and multi-step planning, such as graph algorithms and dynamic programming problems, which often necessitate customized logical deduction. This performance gap aligns with our earlier findings: a notable strength in template-driven tasks but a weakness in complex reasoning.
A.3 Performance across Pass@k Settings
Since we generate the dataset using o4-mini-medium, we study the pass@k performance directly on it. Pass@k metric defines a problem as solved if any of the top-$k$ generated candidates passes all test cases. The results are shown in Table 2.
Table 2: o4-mini (medium) performance for pass@k settings
k 1 2 3 4 5 6 7 8 9 10
Pass Rate (%) 66.1 73.6 77.4 79.8 81.4 82.5 82.8 83.5 83.5 83.5
The pass rate rises from 66.1% at $k=1$ to 83.5% at $k=10$, an improvement of over 20 percentage points. This indicates that repeated attempts significantly enhance performance, suggesting that many problems require multiple sampling to solve. The performance plateaus after $k=8$, implying diminishing returns beyond this point. Even with multiple attempts, the model does not achieve a near-perfect score, illustrating the benchmark has high upper limit and diagnostic value. These results underscore the importance of sampling numerous candidates for difficult tasks and reflect the complexity and variability inherent in the problems.
To analyze the intrinsic difficulty and recoverability of different problem categories, we group algorithm tags into three classes based on the marginal gain in pass rate from $\text{pass}@1$ to $\text{pass}@3$:
  • Significant Improvement ($\geq 20\%$). This category includes tree algorithms, graph algorithms, and mathematical problems. These tasks often admit multiple valid solution paths or implementation strategies; consequently, sampling multiple candidates substantially increases the probability of success.
  • Moderate Improvement ($10\%\sim 20\%$). This group comprises tags such as string algorithms and data structures. These problems typically follow established algorithmic templates, where additional sampling mainly helps mitigate localized implementation errors or off-by-one bugs.
  • Minor Improvement ($\leq 10\%$). This category features dynamic programming and greedy algorithms. The limited benefit of increased sampling suggests that these problems either have a high baseline pass rate or rely primarily on rigorous structural reasoning rather than implementation variability.
The disparity in improvement across categories indicates that multi-candidate evaluation is highly effective for problems with diverse solution pathways, but yields diminishing returns for tasks that require deep logical reasoning or systemic correctness.
Figure 8: Per-tag improvement from pass@1 to pass@3. Tags are grouped by improvement magnitude to illustrate which problem classes benefit most from candidate diversification.
A.4 Alignment with Contamination-Free Code Benchmarks
To verify the validity and robustness of UniCode, we evaluate its alignment with two widely recognized, contamination-free benchmarks: LiveCodeBench and LiveCodeBenchPro. These benchmarks are specifically designed to mitigate the effects of data leakage, providing a reliable gold standard for performance comparison. To ensure experimental integrity, we utilized a consistent ensemble of representative models across all evaluations and conducted a Pearson correlation analysis to quantify the relationship between UniCode and these established metrics.
Figure 9: The alignment between UniCode and established benchmarks (LiveCodeBench and LiveCodeBenchPro). The degree of alignment we achieve, with reference to the absolute value of the correlation coefficient $r$, surpasses the inter-correlation among the established benchmarks.
The experimental results, as visualized in Figure 9, reveal a high degree of statistical consistency: we observe a strong positive correlation with LiveCodeBench ($r>0.9$), indicating that UniCode’s evaluation results are highly congruent with the scoring mechanisms of LiveCodeBench, and a strong negative correlation with LiveCodeBenchPro (r<0.9r<-0.9), which is expected given the differing scoring conventions—while UniCode follows a “higher-is-better” metric, LiveCodeBenchPro employs a ranking-based system where lower numerical values indicate superior performance.
The high absolute correlation values ($|r|>0.9$) across both benchmarks provide empirical evidence that UniCode serves as a reliable and high-fidelity proxy for model performance. This alignment confirms that our benchmark effectively captures the underlying coding capabilities of models in a manner.
A.5 Performance Stability via Recursive Augmentation
To further investigate the depth and resilience of Large Language Models’ (LLMs) reasoning capabilities, we conducted a recursive augmentation experiment. This process introduces cumulative structural and logical shifts, moving the test cases further away from the data distributions potentially encountered during pre-training.
Experimental Setup
We randomly sampled a subset of 132 problems from Variant-L1 and re-applied the UniCode framework to generate 68 Variant-L2 problems. The evaluation spans three progressive levels:
  • Seed: Original problems curated from human-centered benchmarks (e.g., Codeforces).
  • Variant-L1: First-generation variants generated based on the Seed problems.
  • Variant-L2: Second-generation variants generated by applying UniCode’s augmentation logic to Variant-L1.
Findings and Analysis
As illustrated in ,a consistent performance collapse is observed across all tested models. While models exhibit varying degrees of resilience at L1, the transition to L2 leads to an additional average decline of 7.9%.
Interestingly, the performance drop from L1 to L2 is generally less severe than the initial collapse from Level 0 to L1. We posit that the transition from Seed to L1 primarily serves to decouple statistical shortcuts and rote memorization of canonical training data. Once these shortcuts are neutralized at L1, the subsequent move to L2 tests the model’s intrinsic reasoning depth rather than further memory exploitation.
Specifically, the results reveal a divergence in model robustness: Models like o4-mini demonstrate a resilient reasoning core, with a marginal decline of only 2.6% from L1 to L2. This suggests its internal logic remains stable despite increased complexity. In sharp contrast, Gemini-1.5-Flash experiences a substantial drop of 19.4%, indicating that its performance is highly sensitive to even minor structural perturbations once canonical patterns are removed.
Conclusion
This recursive evaluation confirms that current static paradigms severely overestimate model intelligence by conflating memorization with reasoning. UniCode’s ability to generate deeper-level variants provides a more realistic and rigorous ”upper bound” for evaluating the genuine logical capacities of LLMs.
Figure 10: Robustness Decay under Recursive Augmentation. Accuracy of various LLMs on the original seed problems (Seed), first-order variants (Variant-L2), and second-order variants (Variant-L2) generated by UniCode.
A.6 Code-tag Distribution
To systematically evaluate code generation capabilities of large language models (LLMs), we constructed a hierarchical taxonomy that organizes algorithmic knowledge into tags, subtags, and atomic skills. In total, the taxonomy consists of 9 top-level tags, 31 subtags, and 161 skills, covering both fundamental algorithms and advanced techniques. Figure 11 provides a summary of this distribution.
This taxonomy brings several advantages for code generation evaluation. First, it ensures broad algorithmic coverage: the tags span essential paradigms such as graph algorithms, dynamic programming, data structures, and mathematical methods, allowing evaluations to probe diverse coding skills. Second, the inclusion of fine-grained subtags and skills provides diagnostic granularity. Instead of producing only aggregate scores, we can profile model performance across different algorithmic domains, exposing specific strengths (e.g., string hashing, greedy heuristics) and weaknesses (e.g., bitmask dynamic programming, numerical stability). Third, tagged organization supports balanced dataset construction, ensuring that evaluations are not biased toward a narrow set of skills. It also facilitates longitudinal comparisons: since the taxonomy is stable, we can track progress across model iterations and architectures. Finally, many of the listed skills, such as hashing and network flow, are directly relevant to industrial software engineering and competitive programming, thereby improving the real-world applicability of the evaluation.
Figure 11: Distribution of tags, sub-tags, and skills in UniCode dataset. Best viewed when zoomed in.
A.7 Example Problems
Seed Problem 1: Path of Tasty Dishes
Problem Statement: Read problems statements in Mandarin Chinese and Russian. Suraj, the Chief Prankster is back in action now and this time he has stolen the valentine’s day gift given by Ashi (the love of Chef) to the Chef and ran away with it to Byteland.
Byteland is a not a regular place like Chef’s town. The safest way from Chef’s town to Byteland is through the path of tasty dishes. The path is named so because there are magical tasty dishes which appear to the traveler that no one can resist eating. Also, Suraj has added a strong sleep potion to each of the dish on this path to stop anyone from following him.
Knowing the devilish nature of Suraj, Ashi is concerned about the Chef and has asked all of Chef’s town people to help. The distance from Chef’s town to Byteland through the path of tasty dishes is $X$ units. They have the location where the magic dishes are and how many people are required to eat it completely. Anyone who eats a dish would go to a long sleep and won’t be able to continue. They have the information about the tribal clans that live along the path of tasty dishes who can be of real help in this journey.
The journey Chef and his friends can be described as follows: There is a total of $B$ dishes on the path. Each dish is located at distance $x_{i}$ ($x_{i-1}<x_{i}$). To minimize the number of friends Chef has to leave behind, all of them have decided that exactly $y_{i}$ of them will eat the $i^{th}$ dish. Also, there are $C$ tribal chef clans. Clan $i$ is located at distance $p_{i}$ ($p_{i-1}<p_{i}$) with a population of $r_{i}$. If a group of at least $q_{i}$ men approaches them, they will join the forces.
Input Format:
  • The first line contains an integer $T$, the number of test cases.
  • Each test case contains:
    • Line 1: $X$ (distance to Byteland).
    • Line 2: $B$ (number of dishes).
    • Line 3: $B$ pairs of space-separated integers $x_{i},y_{i}$.
    • Line 4: An integer $C$, followed by $C$ space-separated triplets $p_{i},q_{i},r_{i}$.
Output Format:
For each test case, print the minimum size of the group (including Chef) needed to reach Byteland.
Constraints:
  • $1\leq T\leq 10,\quad 1\leq X\leq 10^{9},\quad 1\leq B\leq 10000$
  • Subproblem 1 (25 pts): $C=0$
  • Subproblem 2 (75 pts): $1\leq C\leq 10000$
  • $1\leq x_{i}<X,\quad 1\leq p_{i}<X$ (all positions are distinct)
  • $1\leq y_{i},q_{i},r_{i}\leq 10^{14}$
Seed Problem 2: Baratheon’s Reign
Problem Statement: The Baratheons have been ruling in the Seven Kingdoms for many years. King Joffrey Baratheon commanded to build two monuments. The Baratheons have been ruling for $N$ years. Every year is described by an integer $A_{i}$, the level of prosperity.
You are to pick two historical periods $S_{1},F_{1}]$ and $S_{2},F_{2}]$ with the following rules:
  • No overlap: Two periods shouldn’t have common years.
  • Chronological: The first period must start earlier than the second one.
  • Separation: There must be at least $K$ years between $F_{1}$ and $S_{2}$.
Goal: Maximize the total sum of prosperity levels in the chosen periods.
Input Format:
  • Line 1: $T$ (test cases).
  • Each case:
    • Line 1: $N$ (years) and $K$ (gap).
    • Line 2: $N$ integers $A_{1},A_{2},\dots,A_{N}$.
Output Format:
For each test case, output a single line containing the maximum sum.
Constraints:
  • $1\leq T\leq 5,\quad 2\leq N\leq 10^{5},\quad 0\leq K\leq 10^{5}$
  • $10^{9}\leq A_{i}\leq 10^{9},\quad K+2\leq N$
New Problem: Chef’s Grand Expedition
Problem Statement: Chef must journey in two phases.
Phase 1: Recruitment
  • There are $N$ districts in Chef’s town, labeled $1\dots N$. District $i$ has $A_{i}$ potential volunteers ($A_{i}$ may be negative: a negative value means the district actually shuns the effort).
  • Chef may conduct exactly two recruitment campaigns, each on a contiguous interval of districts $[L,R]$. These two intervals must not overlap, and there must be at least $K$ districts between the end of the first and the start of the second.
  • Chef gathers the sum of $A_{i}$ in each chosen interval. His total recruits $H$ is the sum over both intervals (if that sum is negative, he would of course choose intervals giving non-negative sum).
Phase 2: Expedition
  • The path from Chef’s town to Byteland has $B$ magical “dishes” at strictly increasing distances $x_{i}$. To cross dish $i$, exactly $y_{i}$ members of Chef’s party must stop (and thus be lost to sleep).
  • There are $C$ tribal clans at strictly increasing distances $p_{j}$. Clan $j$ will join Chef’s party and contribute $r_{j}$ people, but only if at the moment Chef arrives at $p_{j}$ his current party size is at least $q_{j}$.
Chef starts Phase 2 with $G_{0}+H$ people, where $G_{0}$ is the size he sets aside before recruitment. As he moves in increasing order of position he encounters dishes and clans. He must ensure that at every dish he has $\geq y_{i}$ people (to send them to sleep) and that after subtracting $y_{i}$, his party remains >0>0. Similarly, at each clan he gains $r_{j}$ if current $\geq q_{j}$.
Goal: Compute the minimal $G_{0}$ such that Chef can complete Phase 2 alive.
Input Format:
  • Line 1: $T$ (number of test cases).
  • For each test case:
    • Line 1: $N$ and $K$.
    • Line 2: $N$ space-separated integers $A_{1},A_{2},\dots,A_{N}$.
    • Line 3: $B$ (number of dishes).
    • Next $B$ lines: Two integers $x_{i},y_{i}$ for the $i$-th dish.
    • Next line: $C$ (number of clans).
    • Next $C$ lines: Three integers $p_{j},q_{j},r_{j}$ for the $j$-th clan.
Output Format:
For each test case, print one integer: the minimum $G_{0}$.
Constraints:
  • $1\leq T\leq 5,\quad 1\leq N,B,C\leq 10^{5},\quad 0\leq K<N$
  • $10^{9}\leq A_{i}\leq 10^{9}$
  • $1\leq x_{i}<x_{i+1}\leq 10^{9},\quad 1\leq y_{i}\leq 10^{14}$
  • $1\leq p_{j}<p_{j+1}\leq 10^{9},\quad 1\leq q_{j},r_{j}\leq 10^{14}$
Appendix B Technical Implementation and Reproducibility
B.1 Test Cases Quality and Ablation Study
We evaluate test suites using two metrics: correctness (accepting valid solutions) and coverage (rejecting invalid ones). An ideal test suite optimizes both axes simultaneously. Let $S_{\text{correct}}$ and $S_{\text{incorrect}}$ be sets of correct and incorrect submissions. A test suite $M$ passes a submission $s$ if $s$ succeeds on all test cases $m_{i}\in M$, denoted $\text{pass}(s,M)=1$. We define:
$\text{Corr@N}=\frac{|\{\,s\in S_{\text{correct}}\mid\text{pass}(s,M)=1\,\}|}{|S_{\text{correct}}|},$ (1)
$\text{Cov@N}=\frac{|\{\,s\in S_{\text{incorrect}}\mid\text{pass}(s,M)=0\,\}|}{|S_{\text{incorrect}}|}.$ (2)
Table 3: Quality evaluation of generated test suites. We compare our full pipeline against the rStar-Coder baseline (Liu et al., 2025) and perform an ablation study to quantify the contribution of each component. Stage 1 (Brute-Force) uses brute-force solvers on small random inputs. MajVote (Unfiltered Solvers) applies a majority vote to all generated solutions without pre-filtering. MajVote (Filtered Solvers) uses only solutions that passed the Stage 1 stress test. Our Full Pipeline integrates all stages and input types. “Validated suites” is the percentage of problems for which a suite passed all validation checks.
Method Input types Correctness Coverage
Problems with
validated suites (%)
rStar-Coder MajVote (Unfiltered) randrand 86.9% 80.2% 94.3%
Ours Stage 1 (Brute-Force) randrand 91.9% 81.5% 98.2%
MajVote (Unfiltered) allall 86.7% 85.2% 93.9%
MajVote (Filtered) allall 93.8% 84.3% 92.8%
Full Pipeline allall 94.5% 86.0% 94.8%
Setup.
We evaluated 80 problems from the Test-Eval dataset (Yang et al., 2025b), each having an average of 200 pass/fail solutions, using test suites of size $N=50$ (see Section B.2 for test suite composition). Note that our approach to majority voting differs from rStar-Coder in its granularity. While rStar-Coder aggregates at the solution level and discards a problem unless a majority of solutions exhibit identical input–output behavior; our aggregation is performed per test case, so disagreement on individual cases does not invalidate the entire problem.
Results and Ablation Analysis.
As shown in Table 3, our full pipeline significantly outperforms the baseline in both correctness (94.5% vs. 86.9%) and coverage (86.0% vs. 80.2%). This performance gain is dissected through a series of ablation experiments:
  • Impact of Stage 1 Filtering: Comparing MajVote (Unfiltered) with MajVote (Filtered) reveals that our brute-force validation on small inputs is pivotal for correctness, yielding a +7.1% absolute improvement (86.7% to 93.8%) by eliminating erroneous solver candidates.
  • Contribution of Adversarial Inputs: The transition from Stage 1 (random inputs) to the Full Pipeline highlights the role of adversarial and corner inputs in enhancing coverage (increasing from 81.5% to 86.0%), proving their efficacy in exposing complex-case failures.
  • Robustness of Aggregation: Despite the increased complexity, our per-test-case adjudication maintains a high “validated suite” rate (94.8%), outperforming rStar-Coder’s rigid solution-level majority vote.
While the automatic generation system can not be error-free, our analysis in App. B.5 confirms that the resulting benchmark remains statistically reliable for evaluating code generation.
B.2 Test-Suite Composition and Parameter Selection
In this section, we provide a detailed justification for the composition of the final test suite $S$, which consists of 50 test cases with a fixed distribution: 20 random ($I_{\text{rand}}$), 20 adversarial ($I_{\text{adv}}$), and 10 corner ($I_{\text{corn}}$) inputs. This configuration was determined through an extensive empirical evaluation aimed at balancing correctness and coverage.
Experimental Setup
We conducted a hyperparameter sweep over multiple test suite compositions, evaluating each configuration on a held-out set of 48 problems with 960 human-crafted solutions. Each configuration was assessed using two key metrics:
  • Correctness: the proportion of valid solutions that pass all test cases.
  • Coverage: the proportion of invalid solutions that are correctly rejected.
Table 4: Representative sweep results across different distributions.
Distribution Correctness (%) Coverage (%)
(5, 5, 0) 97.9 77.0
(10, 10, 5) 95.8 81.4
(20, 20, 10) 94.0 87.5
(30, 30, 20) 91.7 88.7
(50, 50, 20) 91.7 90.0
We observe that smaller test suites (5,5,0) achieve high correctness but suffer from low coverage, failing to detect many faulty solutions. Larger suite such as (30,30,20), improves coverage marginally but at the cost of increased sensitivity to corner test cases and higher computational cost. The configuration (20, 20, 10) is selected as the optimal ”elbow point. It maintains high correctness 94.0% while achieving broad coverage 87.5%, providing a rigorous filtering mechanism without being prohibitively expensive or overly punitive to valid code.
B.3 Analysis of Generator Bias
A potential concern in generative evaluation is the risk of “generator bias”, where a model performs better on problems it generated itself. To address this concern, we performed validation using an alternative, open-source generator deepseek-r1 to generate a new set of 104 problems across 5 distinct tags. We then benchmark 6 models of varying capability levels on this independently generated set and compared the results to their performance on the standard o4-mini-generated UniCode problems and human-curated no data contamination LiveCodeBench [^2] Initial Release: 8/1/2024 to 5/1/2025.
Model Unicode (deepseek-generated) Unicode (o4mini-generated) LiveCodeBench (human-curated, no contamination)
gpt-5 72.5% 67.7%
o4-mini 70.2% 66.9% 74.2%
deepseek-r1 61.6% 56.6% 73.1%
o3-mini 51.0% 57.1% 63.0%
gpt-4.1-mini 41.3% 42.4% 53.2%
gemma-3-27b-it 14.6% 13.0%
Table 5: Pass@1 rates across different problem generators. While absolute scores show fluctuation across the three datasets, indicating differences in generator difficulty or stylistic preferences. The relative model hierarchy remains highly consistent (Pearson $r=0.984$ between deepseek-r1 and o4-mini generated sets).
As indicated in Table 5, the results do not support the presence of significant self-preference bias. The relative ranking of models remains highly consistent across all three datasets, with a Pearson correlation of $r=0.984$ ($p=4\times 10^{-4}$) between model performances on the deepseek-r1-generated and o4-mini-generated problem sets.
Although deepseek-r1 performs markedly better on its own problems than on o4-mini’s (61.6% vs. 55.6%), a similar performance gap persists on human-curated LiveCodeBench. This suggests the difference stems from stylistic preferences inherent to each generator, rather than intentional bias. Therefore, to mitigate any residual stylistic bias and ensure robust evaluation, we will release UniCode-Multi benchmark, a composite benchmark aggregating $100$ problems each from o4-mini, gpt-5, gemini-2.5-pro, Deepseek-r1, and qwen3-235b-a22b models.
B.4 Validity of Generated Problems
To rigorously assess the utility and complexity of our generated benchmark, we conducted an extensive human evaluation involving 113 problems (approx. 23% of the total problem set). This process required over 30 hours of expert labor, as each problem underwent 10–20 minutes of in-depth analysis by competitive programming veterans.
Expert Annotation Protocol
We recruited five independent annotators: senior competitive programmers and algorithm engineers with over 5 years of experience and a Codeforces rating of 2100+ (Master level or above). To ensure objectivity, we employed a blinded rating protocol to evaluate the problems:
  • Validity & Rigor. The evaluation yielded a high Validity rate of 98.2%. Technical analysis of the few invalid cases (1.8%) revealed that they were primarily due to minor output specification ambiguities rather than fundamental logical flaws or insurmountable constraints.
  • Inter-Annotator Agreement. We observed a 92.3% agreement rate among the five experts. This level of consensus, especially given the problems’ complexity, ensures the clarity and formal precision of the generated statements.
Figure 12: Human rating website. Best viewed when zoomed in.
Figure 13: Human rating website. Best viewed when zoomed in.
B.5 Trustworthy Evaluation with Erroneous Tasks
Benchmarks for code generation occasionally contain erroneous items (e.g., unsolvable prompts, mislabeled I/O, flawed tests). This section develops a simple contamination model that quantifies how such items affect reported accuracy, provides bias- and variance-aware confidence bounds, and gives practical recipes to maintain trust in benchmark results.
B.5.1 Setup and Notation
Let each task $i\in\{1,\dots,n\}$ be either reliable ($R_{i}=1$) or unreliable ($R_{i}=0$). Write
$\alpha\equiv\Pr(R_{i}=0)\quad\text{and}\quad 1-\alpha\equiv\Pr(R_{i}=1),$
so $\alpha$ is the contamination rate of the benchmark. Let $p\in[0,1]$ denote the model’s true accuracy on reliable tasks and let $q_{e}\in[0,1]$ denote the effective success probability on unreliable tasks (e.g., a random or spurious pass rate). For pass@$k$, define $p$ and $q_{e}$ analogously as the success probability within $k$ attempts.
For each task, the observed outcome $Y_{i}\in\{0,1\}$ indicates success. The reported accuracy is $\hat{\mu}\equiv\frac{1}{n}\sum_{i=1}^{n}Y_{i}$.
B.5.2 Systematic Bias (Identification and De-biasing)
By the law of total expectation,
$\mathbb{E}[\hat{\mu}]=(1-\alpha)\,p+\alpha\,q_{e}\quad\Longrightarrow\quad\mathrm{Bias}(\hat{\mu};p)=\bigl|\mathbb{E}[\hat{\mu}]-p\bigr|=\alpha\,|q_{e}-p|\leq\alpha.$ (3)
Thus, when $\alpha$ is small, the systematic bias is small in absolute value. If $\alpha$ and $q_{e}$ are known (or fixed by design), an unbiased estimator of $p$ is obtained by de-biasing:
$\tilde{p}=\frac{\hat{\mu}-\alpha q_{e}}{1-\alpha}\qquad\text{(exact if $\alpha,q_{e}$ are known).}$ (4)
When only bounds are available, $q_{e}\in[q_{\min},q_{\max}]$ and $\alpha\in[0,\alpha_{\max}]$, one obtains a conservative identification region for $p$:
$p\in\biggl[\,\frac{\hat{\mu}-\alpha_{\max}q_{\max}}{1-\alpha_{\max}},\ \frac{\hat{\mu}-\alpha_{\min}q_{\min}}{1-\alpha_{\min}}\,\biggr]\cap[0,1].$ (5)
In practice, $q_{\max}$ can be set by a null-model baseline (e.g., trivial solver or random program generator), and $\alpha_{\max}$ by audit sampling.
B.5.3 Random Error (Sampling Variability)
There are two natural regimes for variance, depending on whether the reliable/unreliable split is fixed in advance (e.g., exactly $100(1-\alpha)\%$ reliable) or arises by i.i.d. sampling.
Fixed split (common in controlled curation).
If exactly $(1-\alpha)n$ reliable and $\alpha n$ unreliable tasks are present, [^3] Assuming $n(1-\alpha)$ and $n\alpha$ are integers; otherwise interpret as the nearest integers. then
[…]
Random mixture (i.i.d. contamination).
Marginally $Y_{i}\sim\mathrm{Bernoulli}(\mu)$ with $\mu=(1-\alpha)p+\alpha q_{e}$, hence
$\mathrm{Var}(\hat{\mu})=\frac{\mu(1-\mu)}{n}\qquad\text{with}\quad\mu=(1-\alpha)p+\alpha q_{e}.$ (8)
By concavity of x(1x)x(1-x), (6) $\leq$ (8), so using (8) is conservative when the split is fixed.
Confidence intervals.
Let $z_{0.975}\approx 1.96$. A simple large-sample $95\%$ CI for $\mu$ is
$\hat{\mu}\ \pm\ z_{0.975}\sqrt{\frac{\hat{\mu}(1-\hat{\mu})}{n}}\!,$ (9)
or, more accurately at small $n$, use a Wilson or Agresti–Coull interval for $\mu$. When $\alpha,q_{e}$ are known, a CI for $p$ follows from de-biasing:
$\biggl[\frac{\underline{\mu}-\alpha q_{e}}{1-\alpha},\ \frac{\overline{\mu}-\alpha q_{e}}{1-\alpha}\biggr],$ (10)
where $\underline{\mu},\overline{\mu}]$ is a $95\%$ CI for $\mu$. If only bounds are known ($\alpha\in[0,\alpha_{\max}]$, $q_{e}\in[q_{\min},q_{\max}]$), combine (5) with $\underline{\mu},\overline{\mu}]$ to obtain a conservative CI for $p$:
$\biggl[\frac{\underline{\mu}-\alpha_{\max}q_{\max}}{1-\alpha_{\max}},\ \frac{\overline{\mu}-\alpha_{\min}q_{\min}}{1-\alpha_{\min}}\biggr]\cap[0,1].$ (11)
B.5.4 Total Error Bound
Combining systematic and random components yields a high-probability bound on the absolute estimation error for $p$:
$|\hat{\mu}-p|\ \leq\ \underbrace{\alpha\,|q_{e}-p|}_{\text{systematic bias}\ \leq\alpha}\ +\ \underbrace{z_{0.975}\sqrt{\frac{\mu(1-\mu)}{n}}}_{\text{random error}}\quad\text{(w.h.p.)}.$ (12)
When $n$ is large, the $O(n^{-1/2})$ term vanishes and the total error is controlled by the bias ceiling $\alpha$. If $\alpha$ (and $q_{e}$) are known, report the de-biased estimate (4) with CI (10); this both removes the bias and shrinks the CI.
B.5.5 Stratified (Tag-wise) Contamination
If tasks are grouped into tags $t=1,\dots,T$ with weights $w_{t}$ (sum to $1$), reliable rates $p_{t}$, contamination rates $\alpha_{t}$, and unreliable success $q_{e,t}$, then
$\mu$ $=\sum_{t=1}^{T}w_{t}\bigl((1-\alpha_{t})p_{t}+\alpha_{t}q_{e,t}\bigr),$ (13)
[…]
under a fixed per-tag split. Reporting tag-wise de-biased estimates $\tilde{p}_{t}=(\hat{\mu}_{t}-\alpha_{t}q_{e,t})/(1-\alpha_{t})$ with their CIs, and then aggregating by the $w_{t}$, makes contamination assumptions explicit and auditable.
B.5.6 Numerical Illustration
Take $\alpha=0.06$, $p=0.80$, $q_{e}=0.50$. Then
$\mu=(1-\alpha)p+\alpha q_{e}=0.94\cdot 0.80+0.06\cdot 0.50=0.782,\quad\mathrm{Bias}=|\mu-p|=1.8\%.$
Under a fixed split, the standard error is
$\mathrm{SE}(\hat{\mu})=\sqrt{\frac{0.94\cdot 0.80\cdot 0.20+0.06\cdot 0.50\cdot 0.50}{n}}=\sqrt{\frac{0.1654}{n}}.$
[…]
Table 6: Random error and conservative total error bound (bias ++ half-width) at various $n$ ($\alpha{=}6\%$, p=0.80p{=}0.80, $q_{e}{=}0.50$).
$n$ SE $95\%$ CI half-width Total error bound
500500 $1.82\%$ $3.57\%$ $1.8\%+3.57\%\approx 5.4\%$
5,0005{,}000 $0.575\%$ $1.13\%$ $1.8\%+1.13\%\approx 2.9\%$
10,00010{,}000 $0.407\%$ $0.80\%$ $1.8\%+0.80\%\approx 2.6\%$
As $n$ grows, random error shrinks as $O(n^{-1/2})$; the residual error is then dominated by the (small) bias ceiling $\alpha$.
B.5.7 Practical Safeguards
  • Audit and bound $\alpha$. Spot-check a random subsample to obtain an empirical upper bound $\alpha_{\max}$ with binomial CIs; report $p$ using (11).
  • Calibrate $q_{e}$. Measure $q_{e}$ (or $q_{\max}$) using null models (e.g., trivial programs, permuted I/O) to cap spurious pass rates.
  • De-bias when possible. If $(\alpha,q_{e})$ are fixed by design (e.g., known faulty items), publish the de-biased estimate (4) and its CI (10).
  • Stratify and reweight. Estimate per-tag $(\alpha_{t},q_{e,t})$ and aggregate, reducing sensitivity to heterogeneous contamination.
  • Robust reporting. Alongside $\hat{\mu}$, report (i) de-biased $\tilde{p}$, (ii) contamination-aware CIs, and (iii) sensitivity bands under $(\alpha,q_{e})$ ranges as in (11).
Takeaway.
Even when a benchmark contains a small fraction of erroneous tasks, its reported accuracy remains trustworthy when (i) contamination is explicitly modeled, (ii) bias is de-biased or bounded, and (iii) sampling error is controlled by adequate $n$. In the common regime of small $\alpha$ and large $n$, the total measurement error is tightly bounded and the benchmark reliably reflects true coding performance.
B.6 Complexity Bounds and Time Constraints
To illustrate the relationship between input scale, algorithmic complexity, and execution time, we analyze two approaches to a standard programming problem:
  • Task: Sort an array of $n$ integers and count inversions
  • Input Range: $1\leq n\leq 2\times 10^{7}$
  • Expected Solutions:
    • Optimal: Merge sort ($O(n\log n)$) with inversion counting
    • Suboptimal: Bubble sort ($O(n^{2})$) with brute-force counting
B.6.1 Capacity Analysis
We assume a typical modern computer can perform approximately $10^{8}$ operations per second. We set time limits as 5s for optimized and 50s for brute-force algorithms.
For the optimized algorithm ($T(n)=n\log_{2}n$): we solve $n\log_{2}n\leq 5\times 10^{8}$ to show it can handle input $n=2\times 10^{7}$ in 5 seconds, as the number of operations, $2\times 10^{7}\times 24.2\approx 4.84\times 10^{8}$, stays within the $5\times 10^{8}$ operations limit for 5 seconds at $10^{8}$ operations per second.
For the brute-force algorithm ($T(n)=n^{2}$): $n^{2}\leq 50\times 10^{8}$ yields $n\approx 7\times 10^{4}$ maximum. In contrast, processing input $n=2\times 10^{7}$ would require $4\times 10^{14}$ operations (around 46 days), demonstrating quadratic time growth.
Table 7: Algorithm Capacity Comparison
Metric Optimized ($O(n\log n)$) Brute-force ($O(n^{2})$)
Time Limit 5s 50s
Max $n$ $2\times 10^{7}$ $7\times 10^{4}$
The large difference (approximately 300x) in manageable input sizes ($2\times 10^{7}$ vs $7\times 10^{4}$) explains the stress-driven pipeline: the optimized algorithm verifies efficiency at competition-scale inputs, while the brute-force method allows small-case validation ($n\leq 10^{4}$ in $\leq$ 2s). This setting ensures that the brute-force algorithm has enough time to pass test cases with smaller input sizes, which are usually used to verify basic correctness. This is very useful for debugging and initial testing.
B.7 Integrated Prompts for Test Case Generation
Listing 1: Random Input Generator (CYaRon)
I will provide you with a programming problem description, and your task is to generate standardized test input samples using the CYaRon library.
You need to complete the following steps:
1. Parse the constraints on the input from the problem description, such as the range of input data, specific input constraints, etc.
2. Write a function generate_test_input using the CYaRon library to randomly generate test inputs based on a specified problem size. The function should validate that the parameters fall within the specified constraints. If any parameter is out of range, the function should return None. If the parameters are valid, generate a random test input and return an input string (input_string).
3. Write a function validate_test_input to verify whether the generated test input satisfies the requirements specified in the problem description. This includes checking the input data type and constraints parsed in step 1, such as range and other conditions. The function should take input_string as input and return a boolean (True/False).
Output format (strictly follow)
Part 1: Parse Input Constraints
Specify the input constraints as described in the problem.
Part 2: Code for Test Input Generation
import cyaron as cy #cyaron version: 0.7.0
def generate_test_input():
# set parameters constraints that meet requirements (e.g. 1 <= N <= 300)
# Generate input using CYaRon
input_data = [
[…]
Part 3: Code to Validate Test Input
def validate_test_input(input_string):
# Validation logic
return <boolean>
Note:
- cy.Integer() is not supported; it should be cy.randint.
- use cy.String.random instead of cy.String
- The function generate_test_input() should not accept any parameters. You need to generate the input entirely within the function.
- Generate code following the above format, without starting with ‘‘‘python or similar markers.
Listing 2: Adversarial Input Generator (CYaRon)
I will provide you with a programming problem description, and your task is to generate adversarial test input samples using the CYaRon library.
You need to complete the following steps:
1. Parse the constraints on the input from the problem description, such as the range of input data, specific input constraints, etc.
2. Write a function generate_test_input using the CYaRon library to generate a single adversarial test input designed to challenge boundary conditions or worst-case complexity. The function should internally randomize which adversarial strategy to use, without accepting any parameters. The generated input should still conform to the problem constraints.
3. Write a function validate_test_input to verify whether the generated test input satisfies the requirements specified in the problem description. This includes checking the input data type and constraints parsed in step 1. The function should take input_string as input and return a boolean (True/False).
Output format (strictly follow):
Part 1: Parse Input Constraints
Specify the input constraints as described in the problem.
Part 2: Code for Test Input Generation
import cyaron as cy # cyaron version: 0.7.0
import random
def generate_test_input():
# set parameters constraints that meet requirements (e.g. 1 <= N <= 300)
# Randomly choose one adversarial strategy
strategy = random.choice(["equal_weights", "alternating_large_small", "large_ends"])
if strategy == "equal_weights":
# example: all weights are maximal
N = cy.randint(100000, 100000) # fix to worst-case size
max_weight = cy.randint(109, 109)
k = random.choice([1, N//2, N-1, N])
weights = [max_weight] * N
elif strategy == "alternating_large_small":
N = cy.randint(100000, 100000)
max_weight = cy.randint(109, 109)
k = random.choice([1, N//2, N-1, N])
weights = [max_weight if i%2 else 1 for i in range(N)]
else: # large_ends
N = cy.randint(100000, 100000)
max_weight = cy.randint(109, 109)
k = random.choice([1, N//2, N-1, N])
weights = [max_weight] + [1]*(N-2) + [max_weight]
# Build input string
input_lines = [f"{N} {k}"] + [str(w) for w in weights]
[…]
Part 3: Code to Validate Test Input
def validate_test_input(input_string):
try:
[…]
N_k = lines[0].split()
if len(N_k) != 2:
return False
N, k = map(int, N_k)
if not (1 <= N <= 100000):
return False
if not (1 <= k <= N):
return False
weights = list(map(int, lines[1:]))
if len(weights) != N:
return False
for w in weights:
if not (1 <= w <= 109):
return False
return True
except:
return False
Note:
- generate_test_input() must return a single adversarial input string, not a list.
- Use cy.randint instead of cy.Integer().
- The function should generate adversarial yet valid data fully inside, without parameters.
Listing 3: Direct Test Input Generator
Task:
Generate a challenging test input for the algorithm problem:
{problem_description}
Instructions:
- Focus on edge cases or scenarios that maximize the failure probability in faulty solutions.
- Due to the output length limit, you should generate a small-scale test input that is complete and valid.
- Output the test input directly, not code to generate it.
Output format:
’’’plaintext
{test input}
Think step by step.
B.8 Integrated Prompts for Algorithmic Problem Generation
Listing 4: Prompts for Compositional Variants
You are an expert competitive programmer.
Ill provide you with two programming problems.
If the problems test similar concepts (same-type fusion):
1. Analyze their problem design approaches
2. Create a new challenging problem testing the same concept(s).
If they test different concepts (cross-type fusion):
1. Explore how to combine these concepts
2. Design a new challenging problem that integrates them
you must choose one of the following variation strategies:
1. Sequential Combination: Concepts are loosely chained or simple spliced.
2. Deep Integration (Fusion): Concepts are merged in a non-trivial, deeply interconnected way.
Output format(strictly follow):
## Part 1: Original Problems and Solution Analysis
Step1: [Describe the steps of reasoning]
Step2: xxx
## Part 2: New Problem Description:
New_problem: [Describe the new problem clearly in natural language.]
Input Format: [Specify the input format]
Output Format: [Specify the output format]
## Part 3: Example Test Cases
Input: [Input for test case 1]
Output: [Expected output for test case 1]
Input: [Input for test case 2]
Output: [Expected output for test case 2]
## Part 4: Category
difficulty: [Easy/Medium/Hard]
tags: [tags of new problem, separated by commas.]
variation: [cross-type or same-type fusion; Sequential Combination or Deep Integration.]
Note:
1. Please generate a difficult and original question.
2. The new problem must be rigorous and clearly stated, and include explicit input/output specifications or constraints.
3. Please design questions that have one correct answer; avoid output one possible combination that could have multiple valid answers.
4. Provide two example test cases to demonstrate the new problem.
Listing 5: Prompts for Atomic Variants
You are an expert competitive programmer.
Ill provide you with one programming problem, its solution, and the key concepts they test.
You need to:
1. Analyze its problem design approaches
2. Create a new variation question based on the orignal oneYou must choose **one** of the following variation strategies:
1. **Increase Data Scale:** Significantly increase the constraints, ensuring that an optimized algorithm can still solve the problem within a 5-second time limit.
2. **Rule Transformation (Preserving Core Algorithm):** Change the rules or constraints without affecting the fundamental algorithmic approach (e.g., in a "Climbing Stairs" problem, change steps from {1, 2} to {2, 3}).
Output format(strictly follow):
## Part 1: Original Problems and Solution Analysis
Step1: [Describe the steps of reasoning]
Step2: xxx
## Part 2: New Problem Description:
New_problem: [Describe the new problem clearly in natural language.]
Input Format: [Specify the input format]
Output Format: [Specify the output format]
## Part 3: Example Test Cases
Input: [Input for test case 1]
Output: [Expected output for test case 1]
Input: [Input for test case 2]
Output: [Expected output for test case 2]
## Part 4: Category
difficulty: [Easy/Medium/Hard]
tags: [tags of new problem, separated by commas, referring to the tags of the original problems.]
variation: [The type of variation used: Increase Data Scale or Rule Transformation]
Note:
1. The new problem must be rigorous and clearly stated, and include explicit input/output specifications or constraints.
2. Provide two example test cases to demonstrate the new problem.
3. Select the variation type you are most confident in, aiming to use each of the two types with roughly equal probability over multiple interactions.
4. Please design questions that have one correct answer; avoid output one possible combination that could have multiple valid answers.
B.9 Integrated Prompts for Error Analysis
Listing 6: Prompts for Error Analysis
Task: You are an expert coding problem analyst. You will be given Original Problems, a New Variant Problem, Correct Code, and Wrong Code from different models. Follow the steps below.
#### Step 1: Problem Analysis
* Input: I will provide 1-2 original problems and 1 new variant problem.
* Action: Analyze the knowledge points and key testing points for both the original and new problems.
* Output Format:
> Original Problem 1: [Title]
> - Key Test Point: [Description]
> - Knowledge Point: e.g., Segment Tree, Dynamic Programming
> Original Problem 2: [Title] (if provided)
> - Key Test Point: [Description]
> - Knowledge Point: [Description]
> New Variant Problem: [Title]
> - Key Test Point: [Description]
> - Knowledge Point: [Description]
#### Step 2: Code Error Analysis
* Input: I will provide one correct code solution and several wrong code solutions from specific AI models.
* Action: For each wrong code, analyze its errors by selecting from the predefined Error Taxonomy below. You must identify all errors that the model made.
* Error Taxonomy (Choose from these 4 types):
1. Modeling Error: Selecting an incorrect or unnecessary algorithmic paradigm due to misinterpreting the problem variant (e.g., using greedy instead of DP).
2. Logic/Merge Bug: Wrong conditional logic or incorrect variable usage when merging subproblem results.
3. Indexing/Caching Bug: Incorrect cache sizing, array bounds, or off-by-one errors leading to crashes or incorrect outputs.
4. Complexity Error: Using an algorithm with unnecessarily high time/space complexity (e.g., O(n²) where O(n) suffices).
5. Others: Content is not code, or other implementation errors.
* Output Format for Each Wrong Model:
> Incorrect model: [model name]
> - Error Type: e.g, Modeling Error
> - Reason: [Concise explanation]
> - Erroneous Code Snippet: Code
> Incorrect model: [model name]
> - Error Type: e.g., Logic/Merge Bug, Complexity Error
> - Reason: [Concise explanation]
> - Erroneous Code Snippet: Code
Tips:
Error Types can only be selected from Modeling Error, Logic/Merge Bug, Indexing/Caching Bug, Complexity Error, Others.